# Motivic Homotopy Groups of Spheres and Free Summands of Stably Free Modules

### v2

@@ 3. The sphere spectrum over $k$
For all $s$ and $w$, the following diagram commutes, disregarding the $\cong$” symbols,
$\pi_{s,w}(\mathbb{1})$$\pi_{s}(\mathbb{S})$$\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$$\prod_{p\in\mathcal{P}}\pi_{s}(\mathbb{S}^{\wedge}_{p}),$$\scriptstyle{\cong}\scriptstyle{\cong}$
where the horizontal arrows are given by complex realization and the vertical arrows are completions. When $s\neq 0$, connectivity of the sphere spectrum and Serre’s thesis imply that $\pi_{s}(\mathbb{S})$ is finite so that the right vertical arrow is an isomorphism. When $w\leq\frac{1}{2}s+1$, [10] (for the prime $2$) and [25] (for odd primes) tell us that the lower horizontal map is an isomorphism.
@@
The Beilinson–Soulé conjecture holds that $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))\cong 0$ when $s\geq 1$, independently of $w$. This conjecture is known to hold over number fields, and by [17, Lem. 3.8], it therefore holds for $k=\bar{\mathbb{Q}}$. Therefore we can say that the realization $\pi_{s,w}(\mathbb{1}_{\bar{\mathbb{Q}}})\to\pi_{s}(\mathbb{S})$ is an isomorphism whenever $w\leq\frac{1}{2}s+1$ and $s\geq 1$.
@@
The reason for our preference of the integral cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ to the rational cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$ in Theorem 3.2 is the following observation. Corollary 3.4 holds in full even when (s,w)=(1,1)(s,w)=(-1,-1) and (s,w)=(1,2)(s,w)=(-1,-2) when integer coefficients are used. In both cases, the classical homotopy group is $\pi_{-1}(\mathbb{S})=0$, and the calculations due to [18, Cor. 6.43] and [23, Table 1] give us
$\pi_{-1,-1}(\mathbb{1})\cong{{\mathbf{K}}}^{\mathsf{MW}}_{1}(k)\cong{{\mathbf{K}}}^{\mathsf{M}}_{1}(k)\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(1))$
and
$\pi_{-1,-2}(\mathbb{1})\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(2)).$
Note that in [23, Table 1], all the other terms vanish over an algebraically closed field of characteristic 0, and the spectral sequence concerned collapses at the $\mathrm{E}_{2}$-stage.
We conjecture that the kernel of the realization map is isomorphic to $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ for an algebraically closed field $k$, provided $s\neq 0$.
Remark 3.7.
If $\bar{k}$ is not a subfield of $\mathbb{C}$, but is an algebraically closed field of characteristic 0, a result similar to Corollary 3.4 holds, by virtue of [21]. The realization map may be replaced by a zigzag $\pi_{s,w}(\mathbb{1})\leftarrow\pi_{s,w}(\mathbb{1}_{\bar{\mathbb{Q}}})\to\pi_{s}(\mathbb{S})$, where $\mathbb{1}_{\bar{\mathbb{Q}}}$ denotes the motivic sphere spectrum over $\bar{\mathbb{Q}}$. This zigzag consists of isomorphisms when $0\leq w\leq\frac{1}{2}s+1$. When $w<0$ and $s\not\in\{-1,0\}$, we can say that $\pi_{s,w}(\mathbb{1})$ admits a decomposition as a direct sum of a finite group isomorphic to $\pi_{s}(\mathbb{S})$ and a uniquely divisible group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
@@ 4. Unstable homotopy groups of spheres and Stiefel varieties
We use the results of [3] to extend Corollary 3.4 to $\pi_{s,w}(S^{a,b})$, the unstable motivic homotopy groups of spheres. For the sake of simplicity of exposition, we assume in this section that the ground field $k=\bar{k}$ is embedded in $\mathbb{C}$.
@@
Proposition 4.1 (Asok–Bachmann–Hopkins, [3]).
@@
In the notation of [3], $S^{a,b}\in O(S^{a,b})$. Therefore by [3, Theorem 6.3.4], the fibre $\mathcal{F}$ of the stabilization $S^{a,b}\to QS^{a,b}=\Omega^{\infty}_{\operatorname{P}^{1}}\Sigma^{\infty}_{\operatorname{P}^{1}}S^{a,b}$ lies in $O(S^{p,2b})$ where $p$ is the minimum of 2a12a-1 and a+2b1a+2b-1. Then by [3, Lemma 3.1.19], the motivic space $\mathcal{F}$ is p2b1p-2b-1-$\mathbb{A}^{1}$-connected. Recall that $\mathbb{A}^{1}$-connectivity is measured by coweight, see [18, p. 165]. That is to say
(8) $\bm{\pi}_{a+s,b+w}(S^{a,b})\to\bm{\pi}_{s,w}(\mathbb{1})$
@@
Then the unstable homotopy group $\pi_{d+e}(S^{x+y})$ lies in the stable range and the realization map
@@
is split surjective and the kernel is the subgroup of arbitrarily divisible elements. Additionally
@@
We consider a commutative square in which horizontal arrows are stabilization maps and vertical arrows are $\mathbb{C}$-realization:
@@
The right arrow is split surjective and the kernel is the subgroup of divisible elements, by Corollary 3.4 using inequalities 3 and 4. The two additional claims follow from Corollary 3.4 and Remark 3.5. ∎
@@
To illustrate the result, we provide Figure 1. The figure is supposed to explain what Proposition 4.3 tells us about the complex realization map
$\pi_{d+e,e}(S^{x+y,y})=\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y}).$
@@
We have fixed values for $x,y$ (subject to the condition that $x,y\geq 2$), and indicate regions of interest in the $d,e$-plane.
@@
To the right of the vertical line $d=\min\{2x-2,x+y-2\}$, the motivic Freudenthal suspension theorem does not apply, so our methods break down. Above the diagonal line ey=dx+2e-y=d-x+2, the comparison of the relevant motivic and classical stable homotopy groups of spheres is not understood. When $d+e=x+y$, the stable motivic homotopy group is in the 0-stem, so our results on $\operatorname{Ext}$-completion do not apply. To a lesser extent, this affects the line d+e=x+y1d+e=x+y-1 corresponding to the 1-1-stem, where our understanding of the kernel is worse than elsewhere. To the inequalities of Proposition 4.3 another can be added: when $d<x$, the source group $\pi_{d+e\alpha}(S^{x+y\alpha})$ (corresponding to $\pi_{d-x+(e-y)\alpha}(\mathbb{1})$ of negative coweight) vanishes by a connectivity result of [18].
@@
In Figure 1, we identify four regions in which the comparison map $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ takes on certain characteristics. To the left of the line $d=x$, the source of the map is 0, and we will call this the “0-region”. Second, marked by vertical hatching, there is a closed pentagonal region labelled with the symbol $\cong$”. This is a region where the map is known to be an isomorphism. Third, there is a triangular region marked with crosshatching and labelled “B–S”. Here the map is an isomorphism contingent on the Beilinson–Soulé vanishing conjecture. This region contains its rightmost edge but not the rest of the boundary. Fourth, marked by horizontal hatching, there is a vertical strip with slanted top, containing its boundary, and labelled with “div”. Here the target of comparison map vanishes, and the group $\pi_{d+e\alpha}(S^{x+y\alpha})$ is known to be divisible, and except possibly along the 1-1-stem line, is isomorphic to $\mathrm{H}^{x+y-d-e}(\operatorname{Spec}k;\mathbb{Z}(y-e))$.
$e$ $d$ $d=\min\{2x-2,x+y-2\}$ey=dx+2e-y=d-x+20-stem,  d+e=x+yd+e=x+y1-1-stem, d+e=x+y1d+e=x+y-1weight 1-1,   e=y1e=y-1coweight 0,  $d=x\congdivB$–S
@@
Figure 1. Behaviour of the comparison $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ for fixed $x,y$. Points $(d,e)$ in which $d<x$ are not marked, since $\pi_{d+e\alpha}(S^{x+y\alpha})$ vanishes for dimensional reasons.
@@
The case of the spheres is a base case for an induction that allows us to say something about many Stiefel varieties.
Theorem 4.4.
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+4-r$;
  4. (4)
    $2n\leq e+d$;
The realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
If in addition $n-1\leq e$, then the realization map is an isomorphism.
@@
The commutativity of this diagram is proved in [9, Prop. 5.1]. Map $f_{1}$ is surjective with uniquely divisible kernel by Proposition 4.3: the required inequalities may be easily verified with x=n1x=n-1, $y=n$ and $d+1$ playing the part of $d$. Note that inequalities 14 imply the same inequalities with r1r-1, n1n-1 in place of $r$, $n$. Therefore, the induction hypothesis implies that the map $f_{2}$ is surjective with uniquely divisible kernel. Map $f_{4}$ is surjective with uniquely divisible kernel by Proposition 4.3 again. And map $f_{5}$ is a surjection with uniquely divisible kernel by the induction hypothesis again: inequalities 14 imply the same inequalities with r1r-1, n1n-1 and d1d-1 in place of $r,n$ and $d$. Additionally, the codomains of $f_{2}$ and $f_{4}$ are torsion abelian groups, by the induction hypothesis. It follows that the codomain of $f_{3}$ is also torsion.
It follows from the five-lemma, Proposition A.2, that $f_{3}$ is surjective with uniquely divisible kernel. Lemma 2.5 now assures us that $f_{3}$ is split.
@@
Theorem 4.5.
  1. (1)
    $r\leq n-2$;
  2. (2)
    $d\leq 2n-2r-3$;
  3. (3)
    $e\leq d+3$;
  4. (4)
    $2n\leq e+d$;
  5. (5)
    $n-1\leq e$.
The realization map
$\pi_{d+e\alpha}(V_{r}(\mathbb{A}^{n}))\to\pi_{d+e}(W_{r}(\mathbb{C}^{n}))$
is injective.
@@
Remark 4.6.
If the Beilinson-Soulé conjecture holds for $k$, then the condition $n-1\leq e$ may be replaced by $\min\{n-1,2n-d\}\leq e$ in each of Theorems 4.4 and 4.5. The proofs are unchanged; the result on which they rely, Proposition 4.3, is stronger.

### v3

@@ 3. The sphere spectrum over $k$
For all $s$ and $w$, the following diagram commutes
$\pi_{s,w}(\mathbb{1})$$\pi_{s}(\mathbb{S})$$\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$$\prod_{p\in\mathcal{P}}\pi_{s}(\mathbb{S}^{\wedge}_{p}),$}
where the horizontal arrows are given by complex realization and the vertical arrows are completions. When $s\neq 0$, connectivity of the sphere spectrum and Serre’s thesis imply that $\pi_{s}(\mathbb{S})$ is finite so that the right vertical arrow is an isomorphism. When $w\leq\frac{1}{2}s+1$, [undefj] (for the prime $2$) and [undefah] (for odd primes) tell us that the lower horizontal map is an isomorphism.
@@
Corollary 3.4.
Suppose the Beilinson–Soulé conjecture holds for $k$. Suppose $s$ is a positive integer and $w$ is an integer satisfying $w\leq\frac{1}{2}s+1$. The realization map
$\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$
is an isomorphism. In particular, this map is an isomorphism when $k$ is the field $\bar{\mathbb{Q}}$ of algebraic numbers.
Proof.
The Beilinson–Soulé conjecture asserts that $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))=0$ when $s\geq 1$, independently of $w$. Assuming this, Corollary 3.3 immediately yields the asserted isomorphism.
This Beilinson–Soulé conjecture is known to hold over number fields by and results of Bloch, explained in [undeff, § 1.6]. By [undefu, Lem. 3.8], the conjecture therefore holds for $k=\bar{\mathbb{Q}}$. ∎
3.1. Further remarks on the homotopy groups of the sphere spectrum
In Figure 1, inspired by [undefj, Fig. 1], we attempt a visual presentation of some of what we know about the groups $\pi_{s,w}(\mathbb{1})$. The $(s,w)$-plane is presented, divided into a number of regions by lines and rays. Light-grey dots indicate a group $\pi_{s,w}(\mathbb{1})$ that is not known to vanish for dimensional reasons. Vanishing of the groups above the line $s=w$ was established by Morel [undefw, Thm. 6.1.8]. We proceed around the rest of the diagram clockwise, starting from the top right corner.
The region marked “finite” is of great interest. The groups $\pi_{s,w}(\mathbb{1})$ appearing here are finite abelian groups for which the comparison $\pi_{r,s}(\mathbb{1})\to\pi_{r}(\mathbb{S})$ is not in general an isomorphism. To $2$-primary eyes, this region is further divided in [undefj, Fig. 1] into a cone above $w=\frac{3}{5}s+1$ consisting of $\eta$-periodic classes and a cone below this line which they call “not understood.” Odd-primary behaviour in this region is less mysterious, but still complicated and interesting, being governed by the differentials in the classical Adams–Novikov spectral sequence according to [undefah, Prop. 2.4.8].
In the region we have marked $\pi_{s}(\mathbb{S})$, on or below the line $w=\frac{1}{2}s+1$, strictly to the left of $s=0$, and on or above the line w=1w=-1, the comparison map $\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$ is an isomorphism. In [undeft], Levine proved that the comparison along the entire $s$-axis is an isomorphism. Our proofs ultimately rely on his result, since the isomorphisms $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})\to\pi_{s}(\mathbb{S}^{\wedge}_{p})$ below the line $w=\frac{1}{2}s+1$ are established by reference to this, combined with an observation that there are no nonzero $\tau$-torsion classes in this region.
Below the line w=1w=-1, there is a quadrant labelled “B–S” for “Beilinson–Soulé”. In this region, the comparison $\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$ is surjective, and has $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ as a kernel. Since $s>0$, the Beilinson–Soulé conjecture is that this kernel vanishes.
Finally, in the third quadrant, there is a region bounded by $s=w$ (inclusive) and s=1s=-1 (exclusive) where there is an isomorphism $\pi_{s,w}(\mathbb{1})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$. For what happens when s=1s=-1, see Remark 3.5.
We have lightly shaded the three diagonal lines of coweight 0, 11 and $2$, where superb results have been obtained. We do not interact with these results much in this paper, but we give a brief account here.
The groups of coweight 0 have been determined by Morel, even unstably as in [undefv, Cor. 6.43]. A detailed analysis of the Adams and the Adams–Novikov-spectral sequences was carried out when $c=1$ in [undefy], allowing the calculation of the sheaf structure on $\bm{\pi}_{1+n\alpha}(\mathbb{1})$ over “low-dimensional fields” (see [undefy] for a precise definition), not just the structure of the global sections over algebraically closed fields. This was superseded, however, by slice spectral sequence techniques in [undefae], [undefaf]. For coweights $c=0,1,2$, these papers give calculations of the sheaf $\bm{\pi}_{c+w\alpha}(\mathbb{1})$ over an arbitrary field. We remark, however, that all the calculations mentioned in this paragraph are determinations of homotopy sheaves in terms of the motivic cohomology of the ground field, so that without further work, they do not establish any cases of the Beilinson–Soulé conjecture.
[…]
Figure 1. The groups $\pi_{s,w}(\mathbb{1})$ over $k=\bar{k}$, with special reference to their comparison with $\pi_{s}(\mathbb{S})$ and $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
@@ 3.1. Further remarks on the homotopy groups of the sphere spectrum
We prefer integral cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ to the rational cohomology $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$ in Theorem 1.1 because the theorem holds in full even when (s,w)=(1,1)(s,w)=(-1,-1) and (s,w)=(1,2)(s,w)=(-1,-2), provided integer cohomology is used. The classical homotopy group in question is $\pi_{-1}(\mathbb{S})=0$, and the calculations due to [undefv, Cor. 6.43] and [undefaf, Table 1] give us
$\pi_{-1,-1}(\mathbb{1})\cong{{\mathbf{K}}}^{\mathsf{MW}}_{1}(k)\cong{{\mathbf{K}}}^{\mathsf{M}}_{1}(k)\cong k^{\times}\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(1))$
and
$\pi_{-1,-2}(\mathbb{1})\cong\mathrm{H}^{1}(\operatorname{Spec}k;\mathbb{Z}(2)),$
both of which are divisible groups. Note that in [undefaf, Table 1], all the other terms vanish over an algebraically closed field of characteristic 0, so that their spectral sequence, the slice spectral sequence for the sphere spectrum, collapses at the $\mathrm{E}_{2}$-page.
Conjecture 3.6.
We conjecture that over an algebraically closed field $k$, the completion map $\mathbb{1}\to\bigvee_{p\in\mathcal{P}}\mathbb{1}^{\wedge}_{p}$ and the unit map $\eta:\mathbb{1}\to\mathsf{M}\mathbb{Z}$ induce isomorphisms
$\pi_{s,w}(\mathbb{1})\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))\oplus\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$
for all $s\neq 0$. This is the content of Theorem 1.1 when $s\neq-1$, so only the case of s=1s=-1 remains unproved.
@@
If $\bar{k}$ is not a subfield of $\mathbb{C}$, but is an algebraically closed field of characteristic 0, a result similar to Corollary 3.3 holds, by virtue of [undefad]. The realization map may be replaced by a zigzag $\pi_{s,w}(\mathbb{1})\leftarrow\pi_{s,w}(\mathbb{1}_{\bar{\mathbb{Q}}})\to\pi_{s}(\mathbb{S})$, where $\mathbb{1}_{\bar{\mathbb{Q}}}$ denotes the motivic sphere spectrum over $\bar{\mathbb{Q}}$. This zigzag consists of isomorphisms when $0\leq w\leq\frac{1}{2}s+1$. When $w<0$ and $s\not\in\{-1,0\}$, we can say that $\pi_{s,w}(\mathbb{1})$ admits a decomposition as a direct sum of a finite group isomorphic to $\pi_{s}(\mathbb{S})$ and a uniquely divisible group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
@@ 4. Unstable homotopy groups of spheres and Stiefel varieties
We use the results of [undefb] to extend Corollary 3.3 to $\pi_{s,w}(S^{a,b})$, the unstable motivic homotopy groups of spheres. For the sake of simplicity of exposition, we assume in this section that the ground field $k=\bar{k}$ is embedded in $\mathbb{C}$.
@@
To pass between stable and unstable motivic homotopy groups, we use the following motivic Freudenthal theorem.
Proposition 4.1 (Asok–Bachmann–Hopkins, [undefb]).
@@
In the notation of [undefb], $S^{a,b}\in O(S^{a,b})$. Therefore by [undefb, Theorem 6.3.4], the fibre $\mathcal{F}$ of the stabilization $S^{a,b}\to QS^{a,b}=\Omega^{\infty}_{\operatorname{P}^{1}}\Sigma^{\infty}_{\operatorname{P}^{1}}S^{a,b}$ lies in $O(S^{p,2b})$ where $p$ is the minimum of 2a12a-1 and a+2b1a+2b-1. Then by [undefb, Lemma 3.1.19], the motivic space $\mathcal{F}$ is p2b1p-2b-1-$\mathbb{A}^{1}$-connected. Recall that $\mathbb{A}^{1}$-connectivity is measured by coweight, see [undefv, p. 165]. That is to say
$\bm{\pi}_{a+s,b+w}(S^{a,b})\to\bm{\pi}_{s,w}(\mathbb{1})$
@@
Then the classical unstable homotopy group $\pi_{d+e}(S^{x+y})$ lies in the stable range, the realization map
@@
is split surjective, and the kernel is the subgroup of arbitrarily divisible elements. Additionally
@@
We consider a commutative square in which horizontal arrows are stabilization maps and vertical arrows are realization maps:
@@
The right arrow is split surjective and the kernel is the subgroup of divisible elements, by Corollary 3.3 using inequalities 3 and 4. The two additional claims follow from Corollaries 3.3 and 3.4. ∎
@@
To illustrate the result, we provide Figure 2. The figure shows what Proposition 4.3 tells us about the groups $\pi_{d+e,e}(S^{x+y,y})=\pi_{d+e\alpha}(S^{x+y\alpha})$, and in particular about the realization map
$\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y}).$
@@
We have fixed values for $x,y$ (subject to the condition that $x,y\geq 2$), and indicate regions of interest in the $d,e$-plane. Light grey dots indicate the presence of groups $\pi_{d+e\alpha}(\mathbb{S}^{x+y\alpha})$ that are not forced to vanish for reasons of connectivity.
The figure may be obtained from Figure 1 by changing stem–weight bigrading to stem–coweight bigrading, moving the origin to account for the passage from stable to unstable homotopy groups, and the drawing of a line $d=\min\{2x-2,x+y-2\}$. The right of this line is where the motivic Freudenthal suspension theorem does not apply, so our methods break down. On or to the left of this line, the theorem applies so that the unstable group $\pi_{d+e\alpha}(S^{x+y\alpha})$ agrees with $\pi_{d-x+(e-y)\alpha}(\mathbb{1})=\pi_{d+e-x-y,e-y}(\mathbb{1})$.
Above the diagonal line ey=dx+2e-y=d-x+2, the comparison of the relevant motivic and classical stable homotopy groups of spheres is not understood. This is labelled “finite” because it corresponds to a portion of the region with the same label in Figure 1.
@@
When $d+e=x+y$, the stable motivic homotopy group is in the 0-stem, which is not torsion so our results on completion do not apply. To a lesser extent, this affects the line d+e=x+y1d+e=x+y-1 corresponding to the 1-1-stem, where our understanding of the kernel is worse than elsewhere. To the inequalities of Proposition 4.3 another can be added: when $d<x$, the source group $\pi_{d+e\alpha}(S^{x+y\alpha})$ (corresponding to $\pi_{d-x+(e-y)\alpha}(\mathbb{1})$ of negative coweight) vanishes by a connectivity result of [undefv].
In Figure 2, we identify regions in which the realization map $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ takes on certain characteristics. These correspond to sub-regions of the similarly named regions of Figure 1 under the comparison between stable and unstable homotopy groups.
@@
To the left of the line $d=x$, the source of the realization map is 0, and we do not label this region further. Second, marked by vertical hatching, there is a closed pentagonal region labelled $\pi_{d+e}(S^{x+y})$”. This is a region where the map is known to be an isomorphism. Third, there is a triangular region marked with crosshatching and labelled “B–S”. Here the map is an isomorphism contingent on the Beilinson–Soulé vanishing conjecture. This region contains its rightmost edge but not the rest of the boundary. Fourth, marked by horizontal hatching, there is a vertical strip with slanted top, containing its boundary, and labelled $\mathrm{H}^{x+y-d-e,y-e}$. Here the target of the comparison map vanishes, and the group $\pi_{d+e\alpha}(S^{x+y\alpha})$ is known to be divisible. Except possibly along the 1-1-stem line, it is isomorphic to $\mathrm{H}^{x+y-d-e}(\operatorname{Spec}k;\mathbb{Z}(y-e))$.
[…]
@@
Figure 2. Some facts about $\pi_{d+e\alpha}(S^{x+y\alpha})$ with reference to the comparison $\pi_{d+e\alpha}(S^{x+y\alpha})\to\pi_{d+e}(S^{x+y})$ for fixed $x,y$. Points $(d,e)$ in which $d<x$ are not marked, since $\pi_{d+e\alpha}(S^{x+y\alpha})$ vanishes for dimensional reasons.
@@
The case of the spheres is a base case for an induction that allows us to say something about many Stiefel varieties. This brings us to our next two main theorems, which we restate for the benefit of the reader. See 1.5
@@
The commutativity of this diagram is proved in [undefi, Prop. 5.1]. Map $f_{1}$ is surjective with uniquely divisible kernel by Proposition 4.3: the required inequalities may be easily verified with x=n1x=n-1, $y=n$ and $d+1$ playing the part of $d$. Note that inequalities 14 imply the same inequalities with r1r-1, n1n-1 in place of $r$, $n$. Therefore, the induction hypothesis implies that the map $f_{2}$ is surjective with uniquely divisible kernel. Map $f_{4}$ is surjective with uniquely divisible kernel by Proposition 4.3 again. And map $f_{5}$ is a surjection with uniquely divisible kernel by the induction hypothesis again: inequalities 14 imply the same inequalities with r1r-1, n1n-1 and d1d-1 in place of $r,n$ and $d$. Additionally, the codomains of $f_{2}$ and $f_{4}$ are torsion abelian groups, by the induction hypothesis. It follows that the codomain of $f_{3}$ is also torsion.
It follows from our modified five-lemma, Proposition A.2, that $f_{3}$ is surjective with uniquely divisible kernel. Lemma 2.7 now assures us that $f_{3}$ is split.
@@
See 1.6
@@
Remark $4.$
If the Beilinson–Soulé conjecture holds for $k$, then the condition $n-1\leq e$ may be replaced by $\min\{n-1,2n-d\}\leq e$ in each of Theorems 1.5 and 1.6. The proofs are unchanged; the result on which they rely, Proposition 4.3, is stronger.